OpenStudy (anonymous):

Simplify (y / y^2 - x^2) + (x / y^2 - x^2) Simplify (y / y^2 - x^2) + (x / y^2 - x^2) @Mathematics

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

y+x / y^2- x^2 = y+x / (y-x).(y+x) =1 /y-x

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

your welcome:)